OpenStudy (anonymous):

having touble solving this one 3 (s+3) - 4 = 3 + 3s + 2

OpenStudy (anonymous):

How did you try?

OpenStudy (anonymous):

3s +9 - 4 = 5 + 3s 3s + 5 = 5 + 3s 3s + 5 - 5 = 5 + 3s - 5 3s = 3s got no further

OpenStudy (anonymous):

Yes, what does that mean then

OpenStudy (anonymous):

Do u divide 3 into each side?

OpenStudy (anonymous):

Ok do it. next

OpenStudy (anonymous):

3s/3s = 3s/3s? s = 1?

OpenStudy (anonymous):

no no, Dividing by 3 you get, s=s isnt it?

OpenStudy (anonymous):

yep but doesn't s = 1?

OpenStudy (anonymous):

and you are supposed to find the value of s for which this condition is true, right! So for what real value of s is s=s

OpenStudy (anonymous):

1 = 1

OpenStudy (anonymous):

anymore?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

is 2=2?

OpenStudy (anonymous):

So you got it right!

OpenStudy (anonymous):

come on tell me the answer

OpenStudy (anonymous):

yes